Daosta Inc. uses the FIFO method in its process costing system. The following data concern the operations of the company's first processing department for a recent month.

Work in process, beginning:

Units in process

900

Percent complete with respect to materials

40

%

Percent complete with respect to conversion costs

20

%

Costs in the beginning inventory:

Materials cost

$

530

Conversion cost

$

2108

Units started into production during the month

16,000

Units completed and transferred out

16,000

Costs added to production during the month:

Materials cost

$

32,180

Conversion cost

$

416,512

Work in process, ending:

Units in process

900

Percent complete with respect to materials

50

%

Percent complete with respect to conversion costs

70

%

Required:

Using the FIFO method:

Equivalent Units of Production for Direct materials

  (Do not enter commas in answer for example 5,000 should be entered as 5000)

Equivalent Units of Production for Conversion Costs

  (Do not enter commas in answer for example 5,000 should be entered as 5000)

Cost per Equivalent Unit – Direct Materials

$ (Round to 2 decimal places)

Cost per Equivalent Unit – Conversion Costs

$ (Round to 2 decimal places)

Total value of Ending Work in Process

$.00 (Round to the nearest whole dollar and do not enter dollar signs or commas in answer for example $5,000 should be entered as 5000)

Total value of units transferred out

$.00 (Round to the nearest whole dollar and do not enter dollar signs or commas in answer for example $5,000 should be entered as 5000)
Answer #1

Using the FIFO method:

Equivalent Units of Production for Direct materials

(900*60%+15100+900*50%) = 16090

Equivalent Units of Production for Conversion Costs

(900*80%+15100+900*70%) = 16450

Cost per Equivalent Unit – Direct Materials

32180/16090 = 2

Cost per Equivalent Unit – Conversion Costs

416512/16450 = 25.32

Total value of Ending Work in Process

450*2+630*25.32 = 16852

Total value of units transferred out

(15640*2+15820*25.32+2638) = 434480